OpenStudy (anonymous):

The prop. blade of an airplane is 2.80ft, and rotates @ 2200 r/min. What is the linear velocity of a point on the tip of the blade?

OpenStudy (zpupster):

Use \[v=\omega(r)=\theta/t(r)\] \[\theta = 2\pi rad\] t= 1 rev r = .5(2.8)

OpenStudy (anonymous):

so them telling me the length of the blade holds no grounds?

OpenStudy (wolf1728):

Will have to look up how to calculate linear velocity.

OpenStudy (anonymous):

all I could find was angular velocity , and after about 2 hours on a simple problem hat

OpenStudy (anonymous):

isnt linear velocity the speed of the circumference?

OpenStudy (anonymous):

speed over time

OpenStudy (anonymous):

I know there's a relationship between the two.

OpenStudy (wolf1728):

I found this: The formula for finding linear velocity is v = x / t. The x variable is distance traveled, and t is the time it took to travel the distance x. The v variable is the linear velocity. radius = 1.4 feet If prop rotates at 2,200 rev/min that equals 36.666 rev for every second Each turn of the prop means that the tip has traveled 1.4 *2*PI = 8.7964594301 feet Multiplying this by 36.666 rev per second equals 322.54 feet per second.

OpenStudy (wolf1728):

The distance from the prop center to the tip is half of 2.8 or 1.4 for each turn of the prop the tip must travel 1.4*2*PI feet.

OpenStudy (anonymous):

so we take 2.80 and multiply in 1/2, that gives us half the blade length which is 1.4 ft. is that 1.4x2xpie?

OpenStudy (wolf1728):

which equals 8.7964594301 feet

OpenStudy (anonymous):

now is that ft per second?

OpenStudy (wolf1728):

eventually it is feet per second. At this point we are only determining the distance traveled - which is 8.796 feet

OpenStudy (anonymous):

persistence pays off i found my page

OpenStudy (anonymous):

I'd like to thank ehuman and wolf1728, I couldn't have done it without you!

OpenStudy (anonymous):

wait what am I talking about, im not done here.

OpenStudy (zpupster):

(2200revs/min)(2 pi rad/rev)(1.4ft)=? that is the equation you need

OpenStudy (zpupster):

compare it to what i gave you before